40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 15 14:10:56 2011, Jenne, Update, SUS, Monday SUS update 7x
    Reply  Mon Aug 15 17:36:10 2011, jamie, Update, SUS, Strangeness with ETMY (was: Monday SUS update) 
    Reply  Tue Aug 30 09:51:37 2016, Steve, Update, SUS, the chamber HEPA tents are back ETMYtentHEPA.jpgvented_for_60_days.png
       Reply  Wed Aug 31 23:01:02 2016, gautam, Update, SUS, the chamber HEPA tents are back 
          Reply  Thu Sep 1 15:40:52 2016, Steve, Update, SUS, HEPAs on each chamber door chamberDoorHepaEA.pdf
Message ID: 12458     Entry time: Wed Aug 31 23:01:02 2016     In reply to: 12453     Reply to this: 12461
Author: gautam 
Type: Update 
Category: SUS 
Subject: the chamber HEPA tents are back 

Some more numbers we found while working in/around the chamber today:

Particle counts
Area 0.3 micron count (per cfm) 0.5 micron count (per cfm)
Vacuum chamber (vented) ITMY  ~5000 ~600
HEPA enclosure on ITMY door <100 0
Lab ~100,000 ~6000

These numbers were measured using our particle counter, which has a pump rate of 0.1 cfm, so the numbers above are 10x the numbers shown on the instrument after a measurement to account for this.

Essentially, the chamber is pretty dirty. Peeling the F.C with hard to reach optics like the ITM installed in place is not really feasible, and after peeling the F.C, we are looking at a best case of an additional 1-2 weeks in air to align the IFO, during which the optic is apparently exposed to quite a lot of particulates. In fact, with the high intensity flashlight left on, I actually saw some flecks of dust occassionally floating around inside the chamber while I was working on the optic. But this is just something we have to accept I guess.

ELOG V3.1.3-